OpenStudy (anonymous):

A clock with a radius of 15 cm has an 11 cm minute hand and a 7 cm hour hand. How far apart are the tips of the hands at 3:30?

OpenStudy (anonymous):

there is 12.5 min between hour and minute when 3:30, because when min takes 30 min, hour takes 2.5 min (because we know that hour takes 5 min when min takes 60 min) so, as we know min between hour and min which is 12.5 min we can calculate angle between them, how, like this, 60 min =360 degree ( min takes 360 degree when passes 60 min) so how many degrees pass 1 min= 6 degree 12.5 *6=75 degree between hour and min you need to use cosine law to find out unknown length which is c^2=a^2+b^2-2abcosC

OpenStudy (anonymous):



OpenStudy (anonymous):

put numbers into the formula c^2=11^2+7^2-2*7*11*cos(75) c=11.4 cm